A class of 40 students is to be divided into smaller groups. If each group is to contain 3,4,or 5 people, what is the largest number of groups possible? Explain your answer.
Answer by drk(1908)

40/3 = 13, r=1
40/5 = 8
40/8 = 5
--
40/3 = 12 + 1 group of 4 for 13 groups